OpenStudy (anonymous):

differentiate 3x + square root x^2 -3

OpenStudy (anonymous):

help

OpenStudy (zpupster):

differentiate 3x first

OpenStudy (anonymous):

so its 3

OpenStudy (zpupster):

then use chauin rule u=x^2 -3 d/du(sqrtu)=??

OpenStudy (anonymous):

teach me how to do the chain rule pls?

OpenStudy (zpupster):

well we are saying u = u=x^2 -3 then d/du(sqrtu) we will take this derivative and plug back in u we know the derivative of (sqrtu) is 1/2sqrtuor plugging back u 1/2sqrt(x2-3) now taking d/dx(x^2-3)= 2x we get 2x/2sqrt(x2-3) the 2's cancel we get 3 + x/sqrt(x2-3)

OpenStudy (anonymous):

ah okay still digesting this.. thanks!!! thank you so much =)

OpenStudy (zpupster):

ok good luck!!